OpenStudy (anonymous):

8x-3(2x+4)=2(x-12) what is x?

Parth (parthkohli):

\(x\) is the 24th letter of the English alphabet.

OpenStudy (skullpatrol):

Do you know how to solve for x?

OpenStudy (anonymous):

there is no root for this equation

OpenStudy (skullpatrol):

$$8x-3(2x+4)=2(x-12)$$

OpenStudy (anonymous):

@parthkohli :you deserve 5 medals and a pat on the back!

OpenStudy (anonymous):

\[S = \emptyset\]

OpenStudy (anonymous):

so all real numbers or no solution ??

OpenStudy (skullpatrol):

What do you think?

OpenStudy (anonymous):

mmmmmhh all real numbers

OpenStudy (skullpatrol):

Why?

OpenStudy (anonymous):

no solution

OpenStudy (skullpatrol):

Just giving the answer does not help.

OpenStudy (skullpatrol):

You must know why.

OpenStudy (anonymous):

i presume its 'all real numbers' bc after i distributed and got rid of parts of the equation i got left at 8x+4=8x-12 (and if i eliminate like this) -8x -8x ughhhhh im lost...

OpenStudy (skullpatrol):

Don't get discouraged you are doing fine, now reason it out...

OpenStudy (anonymous):

\[8x-3(2x+4)=2(x-12) <=> 2x-12=2(x-12) <=> -12=-24 \]

OpenStudy (anonymous):

idk how to go from there would i have x+4=x-12?

OpenStudy (anonymous):

thanks @Hamza-Lol but i want to learn how to do it.

OpenStudy (skullpatrol):

What happens when you subtract x from both sides? $$x+4=x-12$$

OpenStudy (anonymous):

it cancels out?

OpenStudy (skullpatrol):

True, and what equation are you left with?

OpenStudy (anonymous):

4=-12 right?

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

which makes it unsolvable

OpenStudy (skullpatrol):

Not quite.

OpenStudy (anonymous):

grr okay so we got 4=-12 either we divide that or since we dont have an x to solve , we cant solve it and i put x=0 :D ,right?

OpenStudy (skullpatrol):

We don't use the word "unsolvable."

OpenStudy (skullpatrol):

Do you know what a solution set is?

OpenStudy (anonymous):

ohh i get you, teehee good one, but that equation reached it peak right? Making it a 'no solution' like O'boy said?

OpenStudy (anonymous):

idk what were we calling it in highschool but it makes no sens

OpenStudy (anonymous):

no explain please!

OpenStudy (skullpatrol):

A solution set is the set of all solutions of an equation. For example if your equation is $$x=4$$ then the solution set of this equation is {4} because if I replace x with 4 I get the true statement $$4=4$$ Do you understand me so far?

OpenStudy (anonymous):

yess!! but we have a false statement

OpenStudy (skullpatrol):

Now in your question the equation is $$8x-3(2x+4)=2(x-12)$$ and this simplifies to the equation $$4=-12$$ which, as you say, is a false statement. Therefore, the solution set of the original equation has no values of x in it that will make it a true statement. Thus the solution set is empty and this is shown by {} or you could say "no solution."

OpenStudy (anonymous):

wow never thought id get it! thank you !!!!!!!

OpenStudy (skullpatrol):

If you are you ready to test you understanding: solve for x $$2x +3= 2x +3$$

OpenStudy (anonymous):

3=2 no solution :F

OpenStudy (anonymous):

oh no wait i read wrong 3=3 meaning its a true statement

OpenStudy (skullpatrol):

Yes.

OpenStudy (anonymous):

yay ! thank you ! i still need to get to used to these types of equations!

OpenStudy (skullpatrol):

What is the solution set?

OpenStudy (anonymous):

no solution ?

OpenStudy (skullpatrol):

Why?

OpenStudy (anonymous):

because x has no value :S

OpenStudy (skullpatrol):

BTW what grade/age level are you?

OpenStudy (anonymous):

tooooooo embarrassed to say!

OpenStudy (skullpatrol):

Ok that's fine... "no solution" means the solution set is empty, that is because no number will make the equation true.

OpenStudy (anonymous):

but ill tell you that my packet is algebra 1a1, meaning pre-algebra :( im not great with math, whats your grade level do you tutor??? or u just go on here to help?

OpenStudy (skullpatrol):

"all real numbers" means the solution set contains all the real numbers and you can write it as {real numbers} this is the answer to what is the solution set of $$2x+3=2x+3$$ because the equation simplifies to $$x=x$$ which is true for all real numbers.

OpenStudy (anonymous):

i get it thanks again!